Ans: 19 mol

please show the work

What amount of benzene (C6H6) is needed to dissolve 2.10 mol I2(s) to form a solution that boils at 83.7 °C? For benzene, the normal boiling point is 80.1 °C and Kb = 2.53 °C/m.
Answer #1

83.7-80.1 = 2.53 * 2.10/ wt of solvent in solution

3.6= 5.313/ wt of solvent in solution

Wt. of benzene = 5.313/3.6

wt. of benzene= 1.47